OpenStudy (anonymous):

h-21 equal -52 please help

Directrix (directrix):

h-21 equal -52 h - 21 = - 52 Add 21 to both sides. h - 21 + 21 = - 52 + 21 h = ? @jmfelix

OpenStudy (anonymous):

thank you awesome guy

Directrix (directrix):

What did you get for h?

OpenStudy (anonymous):

you mean h

Directrix (directrix):

Right.

OpenStudy (anonymous):

42 and 73

Directrix (directrix):

Nope. - 52 + 21 = ?

OpenStudy (anonymous):

oh wait I forgot it was negative

OpenStudy (anonymous):

73 thanks

Directrix (directrix):

Nope. If you had a bill of 52 dollars to pay and you paid $21 on the bill, you would not still owe $73. ---------------- Do this: 52 - 21 = ? and to the ?, place a negative sign out front.

OpenStudy (anonymous):

-31 ?

Directrix (directrix):

Yes.

OpenStudy (anonymous):

yaaaaaay

Directrix (directrix):

Yes but do study up on addition of signed numbers.